OpenStudy (anonymous):

Which of the following is the equation of a line that passes through the points (0,6) and (2,10)?

OpenStudy (anonymous):

. y = -2x - 6 B. y = -2x + 6 C. y = 2x + 6 D. y = 2x - 6

OpenStudy (anonymous):

plug in the values into the equation. Substitute the (x,y) for the x and y in the equations given, and see which equation proves true for both set of points.